OpenStudy (anonymous):

can someone show me how to find the product (x^2+9)(x^3+6x+3)

OpenStudy (anonymous):

multiply all of \(x^3+6x+3\) by \(x^2\) and then multiply it all by 9 then combine like terms

OpenStudy (anonymous):

x^5+9*9? I am lost

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wow that came out garbled

OpenStudy (anonymous):

\(x^2+9\) has two terms, \(x^3+6x+3\) has three terms so you are going to have to do \(2\times 3=6\) multiplications

OpenStudy (anonymous):

it doesn't matter what order you do them in, but it is easiest to keep track of you multiply first everything in the \(x^3+6x+3\) by \(x^2\) and then multiply all of \(x^3+6x+3\) by 9

OpenStudy (anonymous):

so we start with \[(x^2+9)(x^3+6x+3)=x^2\times x^3+x^2\times 6x+x^2\times 3+9\times x^3+9\times 6x+9\times 3\]

OpenStudy (anonymous):

i am following what you are saying so far

OpenStudy (anonymous):

you usually skip that step, the one i wrote above, and go right to \[x^5+6x^3+3x^2+9x^3+54x+27\]

OpenStudy (anonymous):

then the only like terms to combine are \(6x^3+9x^3=15x^3\) so your "final answer" is \[x^5+15x^3+3x^2+54x+27\]

OpenStudy (anonymous):

I think one of the problems i have is ( ) * ( ) = multiplications but when you are + earlier got confusing when betweent he brackets the first bracket had x^2 second bracket had the + sign how do you know when to add vs using the multiplication sign? does that make sense?

OpenStudy (anonymous):

think of it this way suppose you wanted to multiply \(12\times 213\) you would have to do 6 multiplications, and then at the end some additions. this is like multiplying \((x+2)(2x^2+x+3)\) where \(x\) is ten

OpenStudy (anonymous):

first you would write 213 12 and then you would multiply everything first by 2, then by 1 and then add them up

OpenStudy (anonymous):

got that part

OpenStudy (anonymous):

thanks